OpenStudy (anonymous):

Integral of sin(2x)/cos(x)?

OpenStudy (amistre64):

what is another way of writing sin(2x)?

OpenStudy (amistre64):

this is the reason you learn trigonometry/pre calculus ....

OpenStudy (anonymous):

Wait, does this use the double angle formula?

OpenStudy (amistre64):

it might ....

OpenStudy (anonymous):

you can always use that

OpenStudy (anonymous):

is there a simpler way?

OpenStudy (anonymous):

or a different way altogether

OpenStudy (amistre64):

trig is never simple

OpenStudy (anonymous):

I'm just curious as to the options I have.

OpenStudy (amistre64):

using the wolf lol

OpenStudy (anonymous):

use an abacus

OpenStudy (amistre64):

sin(a+b) = sin(a)cos(b) + sin(b)cos(a) if a=b sin(x+x) = sin(x)cos(x) + sin(x)cos(x) sin(2x) = 2sin(x)cos(x)

OpenStudy (anonymous):

thank you, that's helpful

OpenStudy (anonymous):

I got it.